OpenStudy (anonymous):

Identify each equation as exponential decay or exponential decay. 1. Is this exponential growth or decay?

OpenStudy (anonymous):

y=112(1.320

OpenStudy (anonymous):

@geekfromthefutur @jim_thompson5910

OpenStudy (anonymous):

is y=112(1.320) the complete question? is there an exponent?

OpenStudy (anonymous):

yes its the question @090909090909 no there is an x at the end where the exponent should be but it's not a number

OpenStudy (anonymous):

so it is \[y = 112(1.320)^x\]

OpenStudy (anonymous):

................yes

OpenStudy (anonymous):

to tell if it is growth or decay, look at the number in the parentheses that is being raised to x.

OpenStudy (anonymous):

if that number is between 0 and 1 it is decay, if it is greater than 1 it is growth. so what would it be?

OpenStudy (anonymous):

growth

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5. A city finds its residents are moving to the suburbs. Its population is decreasing by 4% per year. If the initial population of the city was 100,000, what will its population be in 10 years (round answer to the nearest person)?